OpenStudy (anonymous):

Find the amount of time required for an amount to double at the given rate if the interest is compounded continuously. Problem---> 5%

OpenStudy (tkhunny):

\(e^{0.05\cdot t} = 2\) - Solve for t.

OpenStudy (anonymous):

can you explain ? @tkhunny

OpenStudy (tkhunny):

I already did. Have you used logarithms? That's all you need.

OpenStudy (anonymous):

yes , but how would i solve for t?

OpenStudy (anonymous):

@zaynahf

OpenStudy (tkhunny):

Seriously, use logarithms.

OpenStudy (tkhunny):

Base 'e' logarithms? Natural logs? Let's see what you get.

OpenStudy (anonymous):

plug in e^0.05 into my calculator ?

OpenStudy (anonymous):

@zaynahf

OpenStudy (tkhunny):

You did say that you had used logarithms. It appears this is not the case. If: \(e^{0.05t} = 2\) Then: \(\ln\left(e^{0.05t}\right) = \ln(2)\) Further: \(0.05\cdot t\cdot\ln(e) = \ln(2)\) Finally: \(0.05\cdot t = \ln(2)\)

OpenStudy (anonymous):

Then divide both sides by 0.05 ?

OpenStudy (anonymous):

=13.863?

OpenStudy (tkhunny):

That's it. good work. Let's try another one so I can see you deal with the logarithms.

OpenStudy (anonymous):

okay.